a car is moving on ground with speed 3c/5 and a ball thrown from the car with velocity c/2 in upward directiom what is the velocity of ball with respect to ground.
1
Expert's answer
2020-09-09T10:22:01-0400

To calculate the velocity of ball with respect to the ground we need to add velocity of the car and relative velocity of ball in respect to the car (but via relativistic velocity-addition formula):

"u = \frac{v+u'}{1+ \frac{vu'}{c^2}} = \frac{3c/5+c/2}{1+ \frac{3}{10}}= \frac{11}{10} \cdot \frac{10}{13} c = \frac{11}{13} c"